### v2

@@ 1.2 Previous Work
A versatile approach for a robust PD solution is arguably to operate at the block layer, whereby the scheme exposes a block device interface (a common abstraction layer used by many operating systems to represent storage devices as arrays of fixed-size data blocks), providing just a $\mathtt{bRead}$ and a $\mathtt{bWrite}$ method. This is the approach used by solutions like TrueCrypt, and also by Shufflecake; in the remainder of this work, we will only focus on block-layer solutions. In this framework, a single underlying disk or device is formatted to host one or more volumes (logical storage units, usually represented as virtual block devices), each encrypted with a different (usually password-derived) symmetric key. When confronted with the adversary, the user surrenders the passwords to some of these volumes, called decoy volumes (because in some use cases they might contain deceptively innocent data): PD security guarantees that, even after these passwords have been given up, inspection of the disk by the adversary still yields no clue whatsoever hinting at the presence of some further, hidden volumes. The intuition behind the formal definition of these guarantees, is that the adversary cannot distinguish between the case where all the passwords have been surrendered, and the case where there is still undisclosed secret information left.
@@ 1.3 Limitations of Existing Solutions
So far, the landscape of available PD solutions presents many gaps, both in usability and in security, a fact also hinted at by the relatively scarce adoption of such solutions. By far the most widespread today is VeraCrypt, which comes with many limitations. WoORAM-based techniques have been studied in the last few years as promising alternatives to address TrueCrypt and VeraCrypt’s security issues. However, it is important to stress that even the most performant WoORAM-based schemes are still very slow or wasteful. To put this in perspective: HiVE has a slowdown of roughly 200x I/O throughput and wastes 50% of the disk space, while some recent constructions such as DetWoOram [39] reach a slowdown of “only” 5x but at the cost of wasting 75% of the disk space. This leaves us with the dilemma of either choosing single-snapshot, efficient solutions with limited security, or WoORAM solutions with unacceptable performance loss and arguably stronger security.
Moreover, WoORAMs solutions themselves might not be bulletproof. In fact, we believe that the idea that $\mathtt{read}$ requests do not change the underlying state of the physical device is a somewhat strong assumption, and hard to justify with modern, complex SSDs that might, for example, cache $\mathtt{read}$ requests in some undocumented memory area of the firmware, or store read data on an ad-hoc buffer to improve performance, etc.
@@ 1.4 Our Contribution
    It offers a virtually unlimited number of hidden volumes per-device, arranged hierarchically: for the user it is sufficient to unlock one volume, and all the “less secure” volumes will be unlocked automatically. This improves user experience and operational security, as we will see in Section 4.2.
@@
  3. 3.
@@
  4. 4.
    Unlike WoORAM-based solutions, Shufflecake is extremely fast (achieving only a minor slowdown compared to a bare, non-PD system) and wastes less than 1% of the disk space.
Moreover, Shufflecake not only achieves (provable) single-snapshot security, but also implements features that could make possible in the future to achieve a form of “operational” (i.e., weak) multi-snapshot security. These features are Shufflecake’s hierarchical design and atomic block-rerandomisation, which are not available in tools such as TrueCrypt. We discuss this in Section 6.
@@ 2.1 Cryptographic Primitives
Hash functions are designed to be very fast and efficient in term of required computational resources. This might actually be an undesirable property when using the function to store images of user-chosen passwords, because it allows for faster adversarial brute-force. In these cases, a key derivation function (KDF), should be used. KDFs are functionally similar to hash functions, but are designed in such a way to be uniformly expensive to compute on a broad range of computing devices, for example by requiring not only many CPU cycles but also large amount of memory and low latency. Typical KDFs for cryptographic use are Argon2id [5] and Scrypt [34].
@@
If, in addition to privacy, authenticity of a message is also required, then SKEs are not enough, and authenticated encryption (AE) schemes must be used instead. An AE scheme works in a similar way to a SKE, but decryption of a given ciphertext fails if the secret key used for decryption is not the same one used to encrypt the original plaintext. When this happens we write that the decryption procedure returns $\bot$. This allows to check that a ciphertext has not been altered or replaced by a malicious adversary, thereby granting authenticity and integrity of the message. A typical way to implement AE is to append a message authentication code (MAC) to a ciphertext. A MAC is a random-looking bitstring, for example computed through the encrypt-then-MAC procedure with a hash function on a combination of ciphertext and secret key. MACs are useful, among other properties, to check whether a provided key is the correct one to decrypt a ciphertext (without having to actually decrypt the ciphertext first).
@@
As a building block for SKEs, block ciphers are widely used. These are algorithms that typically offer two different interfaces, one for encryption and one for decryption. In encryption mode, they take as input a block of plaintext of fixed bitsize $S_{\mathsf{B}}$ (the block size) and an encryption key of $K$ bits (the key size) and return a block of ciphertext, also of size $S_{\mathsf{B}}$. This mapping is undone in decryption mode, provided the same key is used. In other words, block ciphers implement a subset of size at most $2^{K}$ of the space of all possible $(2^{S}_{\mathsf{B}})!$ permutations (and their inverses) over $S_{\mathsf{B}}$-bit strings. One of the most widely used block ciphers are those from the AES family [20], identified by AES-$K$ (with a block size of 128 bits, and a keysize $K$ of 128, 192, or 256 bits).
@@ 2.2 Full Disk Encryption
FDE can be implemented using hardware-based or software-based solutions. Hardware-based FDE is typically performed by a dedicated encryption chip, while software-based FDE is achieved using encryption software that runs at the operating system level. Implementation is usually done using standard block ciphers contructions and modes of operations like AES-CTR-256. Some examples of software-based FDE solutions include BitLocker for Windows, FileVault 2 for macOS, and LUKS for Linux [30, 15]. All these implementations have typically only negligible impact on performance compared to a non-FDE system, also thanks to the widespread presence of dedicated CPU instructions to speed up AES computation on personal devices such as smartphones and laptops.
@@ 2.3.2 PD security
Let us first clarify that some constraint is necessary in order to have any hope in the PD game against the adversary. Otherwise, the adversary could submit $O_{0}$ and $O_{1}$ containing completely different (logical) $\mathtt{write}$ accesses to the decoy volumes $V_{0},\ldots,V_{N-2}$, and there would be no way of making the two outcomes indistinguishable, since the adversary holds the passwords to those volumes, so it could trivially verify which of the two patterns was executed. This suggests the need for a minimal rule, stating that, whether $O_{0}$ or $O_{1}$ is executed, the resulting logical contents of the decoy volumes $V_{0},\ldots,V_{N-2}$ must be the same. From the user’s perspective, this basic requirement means that we do not try to disguise the $\mathtt{write}$s to decoy volumes as something else, both because we do not need to, and because there would be no way of doing it even if we wanted to.
Furthermore, we notice that many PD solutions (including those based on WoORAMs) treat $\mathtt{write}$ requests in a completely different way than $\mathtt{read}$ requests, i.e. a $\mathtt{write}$ request could trigger allocation or reshuffling of a certain volume sector. This might happen without breaking the minimal rule described above: if an adversary first $\mathtt{read}$s at a previously unallocated position $\beta_{\mathsf{L}}$, obtaining data $d$, and then $\mathtt{write}$s the same data $d$ at the same location $\beta_{\mathsf{L}}$, this might cause a change of state in the instance without changing its logical content (as mandated by the minimal rule above). This, in turn, would enable an attack where the adversary merely checks whether this state change happened or not. However, we must consider this attack trivial: detecting this kind of state change on a decoy volume would actually not compromise security. Therefore, in order to capture this concept, we also demand that the set of all blocks that are touched by $\mathtt{write}$ requests be the same for both $O_{0}$ and $O_{1}$, for all volumes $V_{0},\ldots V_{N-2}$. We stress that this extra constraint is also very minimal and does not weaken the security guarantees offered in practice; in fact, most schemes demand even stricter constraints, for example by demanding that $O_{0}$ and $O_{1}$ are of the same length, which we do not require here.
@@ 3.1 Design
One big limitation of TrueCrypt is that it can only support a hidden volume if the outer volume is formatted as a FAT filesystem. This is because we need the empty space left by the decoy volume to be contiguous. Most modern filesystems (like ext4 and NTFS) “jump back and forth” on the disk as they are written with data, leaving lots of empty blocks, or “holes”, in the middle. Instead, the FAT filesystem is special, in that it grows incrementally and, up to the physiological holes created by deleted files, occupies all the space up until its last utilised block. This way, one can have the hidden volume start at a certain offset, after the end of the decoy volume, and then follow data allocation linearly. TrueCrypt automatically computes a convenient starting position for the hidden volume (leaving some leeway buffer after the end of the standard volume), and places it among the metadata of the hidden volume’s header. Whereas the hidden volume is assigned a logical size that follows the physical space actually allocated to it, the standard volume is not resized, and keeps logically mapping onto the whole disk. This is crucial in order to not defy deniability: if we resized the standard volume, this information (which leaks the existence of a hidden volume) would be written in the metadata of its file system, which is inspected by the adversary.
@@ 3.2 Operational Model
As already mentioned, once the hidden volume is created, its starting position is final, and it “freezes” the end of the decoy standard volume, limiting the maximum size that it will ever be able to attain. Since the standard volume cannot be resized to accommodate for the hidden volume, and instead keeps mapping onto the whole disk, it is up to the user to not let it grow too much and overwrite the hidden volume. This is achieved both by frequent de-fragmentation of the FAT filesystem within the standard volume, and by actually not writing too much data into it. If one wants to be absolutely safe about data corruption within the hidden volume, the recommendation for the user is to never unlock only the standard volume for daily use (except if under coercion), but to always unlock either 1) the hidden volume only, or 2) both hidden and standard volumes, but keeping the latter in read-only mode.
@@ 3.3 Security
In TrueCrypt, once the user surrenders the password of the decoy volume and lets the adversary decrypt it, the only part of the disk contents that remains to be “interpreted” by the adversary is the non-decrypted space after the end of the decoy FAT file system. However, whether this space is actually empty or whether it contains a hidden volume, it will be filled with random bytes that are not readable with the decoy password alone. Therefore, even if a hidden volume is present, the user can plausibly claim that the remaining space is empty and filled with random bytes: the adversary has no way to disprove this claim, or even to question its likelihood, based on the observed disk content.
@@ 3.4 Other Limitations
First of all, as discussed, the standard volume must be formatted as a FAT filesystem if a hidden volume is desired. However, FAT is now outdated: it used to be very widespread, but today there is little plausible reason to use it anymore. Therefore, the mere fact that we are using FAT raises a red flag to the adversary.
@@
Another big limitation is the fact that only one hidden volume within each standard volume is supported. This is a problem, because the adversary might reasonably suspect that TrueCrypt is in use exactly to hide something through its PD feature: if we only meant to encrypt a volume, we would reasonably use a more widespread and supported solution like LUKS or BitLocker. A user could claim that they prefer using independent, niche open-source software to secure their data, and it would be relatively credible, but the safest course of action when designing a PD scheme is assuming that the adversary might not believe this claim, and ask for a second password.
The short answer to this problem is: a robust TrueCrypt-like PD solution should allow simultaneous access to more than two layers of volumes. That way, we can create a series of volumes with increasingly “private” contents (that could well be all decoys), so as to reveal more than one password to the adversary and convince them, based on the resulting decrypted contents, that we have effectively given up what we were hiding, while in fact we are still holding the password to one more top-secret volume whose existence they have no more reason to suspect.
4 The Shufflecake Scheme
In this section we present our Shufflecake scheme, we explain its way of operation, and we provide a security analysis. We do not provide concrete parameters in this section, as we want to present a generic framework that can be adapted to various implementations, but we provide our own choice of parameters in Section 5. In particular, in this section we will treat security as asymptotic, without targeting a concrete bit-security level. We assume a target storage block device of $S_{\mathsf{D}}$ blocks, each of them of size $S_{\mathsf{B}}$ bits, and a maximum of $N_{\mathsf{MAX}}$ supported volumes per device.
@@ 4.1 Design
At its core, Shufflecake is a block indirection layer on top of an encryption layer. Our indirection layer realizes a mechanism which is already a strong improvement over TrueCrypt, since it fixes two of its crucial limitations: it allows multiple volumes, and it is filesystem-independent. Data decryption keys for every volume are derived by a password and other header-derived randomness. Furthermore, the decrypted payload of the header of volume $V_{i}$ with $i>0$ also contains a copy of the header decryption key for volume $V_{i-1}$. This allows to recursively open all volumes present in a device by using a single password, which in turn improves security and user experience, as we will see.
@@
Figure 3: Shufflecake’s disk layout overview.
@@
Figure 4: Shufflecake’s data section layout and slicing.
We distinguish between logical slices, of size $S_{\mathsf{L}}$ blocks, and physical slices, of size $S_{\mathsf{P}}=S_{\mathsf{L}}+\Delta_{\mathsf{S}}$, where $\Delta_{\mathsf{S}}$ blocks are used to store the encryption IVs for that slice. Logical slices store the data of their respective volumes, while physical slices are used to reserve space on the disk to allocate the encrypted logical slice. A physical slice can either be unallocated (i.e., unclaimed by any volume), or mapped to a single logical slice belonging to some volume $V_{i}$. In the latter case, the mapping from the logical slice index (LSI) $\sigma$ of volume $V_{i}$ to the device’s physical slice index (PSI) $\psi$ is given by a correspondence $\mathsf{SliceMap}_{i}:\sigma\mapsto\psi$ which is just a lookup of $V_{i}$’s slice map (basically, a per-volume array). I/O operations to a logical block address $\beta_{\mathsf{L}}$ of volume $V_{i}$ are performed through the two interfaces $\mathtt{SflcRead}$ and $\mathtt{SflcWrite}$. We will describe later these two interfaces, as well as the structure of the slice map.
@@
Figure 5: Shufflecake’s Device Master Block (DMB).
@@
Figure 6: Shufflecake’s Volume Master Block (VMB).
@@
We now describe more in detail how Shufflecake operates and explain the rationale behind certain design choices. We start by explaining how we use encryption to secure data on-disk. Then we look at the indirection layer between physical data on disk and data on the logical volumes, i.e., the mechanism which creates the correspondence between physical and logical slices. We explain how this correspondence is updated when new logical slices are needed as data is written on the logical volumes. Then, we explain how Shufflecake handles the case of allocation errors, i.e., when a physical slice is mistakenly assigned to more than one volume. Finally, we detail how the instantiation procedure works.
@@
In our case, however, we stick to explicitly random IVs because we want to keep the future possibility to extend Shufflecake with some degree of multi-snapshot security, requiring us to re-encrypt some blocks with a different IV while leaving their content unchanged. For this reason, we use CTR mode instead, and the IV of a block is refreshed at each $\mathtt{SflcWrite}$ for that block, to avoid IV-reuse attacks. This means that all these IVs are stored on-disk.
@@
Concretely, we keep an in-memory, per-device array $\mathsf{prmslices}$ of PSIs of size at most $\mathsf{mxslc}=\lceil\frac{S_{\mathsf{D}}-S_{\mathsf{H}}}{S_{\mathsf{P}}}\rceil$ (maximum possible number of slices), and an array of the same size $\mathsf{ofld}$, an “occupation bitfield” telling us which physical slices are occupied (the PSIs are the indexes of the array). Initially, at device instantiation, $\mathsf{prmslices}$ is initialized as $\mathsf{prmslices}[i]:=i$ and then permuted using the efficient Fisher-Yates algorithm [13]. The bitfield $\mathsf{ofld}$ is initialized with all $\mathsf{free}$ values. When a volume is $\mathtt{open}$ed, and we discover the physical slices it maps to, we mark them as $\mathsf{occupied}$ in $\mathsf{ofld}$. The slice allocation algorithm then simply works by repeatedly taking the next element from the pre-shuffled array of PSIs until the bitfield tells us it is free (at which point we take it and mark it as occupied). A stateful occupation counter $\mathsf{octr}$ is kept to facilitate this task, initially set to 00, and then increased up to the first element of $\mathsf{prmslices}$ marked as $\mathsf{free}$ (i.e., the first $\mathsf{octr}$ elements of $\mathsf{prmslices}$ are guaranteed to be $\mathsf{occupied}$ in $\mathsf{ofld}$). This way, when a new slice is required, it is immediate to advance to the next available one. We also need to update $V_{i}$’s slice map $\mathsf{SliceMap}_{i}$ before returning (this is the only change that will eventually persist, encrypted, on disk).

### v3

@@ Abstract
[…]
*                                             A 15-page abstract of this work appears (with the same title) in the proceedings of the ACM Conference on Computer and Communications Security (CCS) 2023. This is the authors’ full version. This document supersedes any previous versions. Original Version: 2023-10-06.      This Revision: 2024-08-23. Reference Software Implementation Version: v0.4.5 (‘Legacy’ scheme only)
@@ 1.2 Previous Work
A versatile approach for a robust PD solution is arguably to operate at the block layer, whereby the scheme exposes a block device interface (a common abstraction layer used by many operating systems to represent storage devices as arrays of fixed-size data blocks), providing just a $\mathtt{bRead}$ and a $\mathtt{bWrite}$ method. This is the approach used by solutions like TrueCrypt, and also by Shufflecake; in the remainder of this work, we will only focus on block-layer solutions. In this framework, a single underlying disk or device is formatted to host one or more volumes (logical storage units, usually represented as virtual block devices), each encrypted with a different (usually password-derived) symmetric key. When confronted with the adversary, the user surrenders the passwords to some of these volumes, called decoy volumes (because in some use cases they might contain deceptively innocent data): PD security guarantees that, even after these passwords have been given up, inspection of the disk by the adversary still yields no clue whatsoever hinting at the presence of some further, hidden volumes. The intuition behind the formal definition of these guarantees, is that the adversary cannot distinguish between the case where all the passwords have been surrendered, and the case where there are still undisclosed ones.
@@ 1.3 Limitations of Existing Solutions
So far, the landscape of available PD solutions presents many gaps, both in usability and in security, a fact also hinted at by the relatively scarce adoption of such solutions. By far the most widespread today is VeraCrypt, which comes with many limitations. woORAM-based techniques have been studied in the last few years as promising alternatives to address TrueCrypt and VeraCrypt’s security issues. However, it is important to stress that even the most performant woORAM-based schemes are still very slow or wasteful. To put this in perspective: compared to a baseline, HiVE has a slowdown of roughly 200x I/O throughput and wastes 50% of the disk space, while some recent constructions such as DetWoORAM [39] reach a slowdown of “only” 5x but at the cost of wasting 75% of the disk space. This leaves us with the dilemma of either choosing single-snapshot, efficient solutions with limited security, or woORAM solutions with unacceptable performance loss and possibly stronger security.
But woORAMs solutions themselves might not be bulletproof. In fact, we believe that the idea that $\mathtt{read}$ requests do not change the underlying state of the physical device is a somewhat strong assumption, and hard to justify with modern, complex SSDs that might, for example, cache $\mathtt{read}$ requests in some undocumented memory area of the firmware, or store read data on an ad-hoc buffer to improve performance through closed-source optimizations, etc.
@@ 1.4 Our Contribution
    It offers a virtually unlimited number of hidden volumes per-device, and such number is always hidden to the adversary.
@@
    Volumes are arranged hierarchically: for the user it is sufficient to unlock one of them, and all the “less secure” ones will be unlocked automatically.
  3. 3.
@@
  4. 4.
    Unlike TrueCrypt, all volumes can be read and written concurrently without integrity pitfalls or security degradation.
  5. 5.
@@
  6. 6.
    Unlike woORAM-based solutions, Shufflecake is extremely fast (with only a minor slowdown compared to a bare, non-PD system) and wastes less than 1% of the disk space.
This document describes the ‘Legacy’ version of Shufflecake, first introduced in 2022. This version of Shufflecake not only achieves (provable) single-snapshot security, but also implements features that could make possible in the future to achieve a form of “operational” (i.e., weak) multi-snapshot security. These features are Shufflecake’s hierarchical design and atomic block-rerandomisation, which are not available in tools such as TrueCrypt. We discuss this in Section 6.
@@ 2.1 Cryptographic Primitives
Hash functions are designed to be very fast and efficient in term of required computational resources. This might actually be an undesirable property when using the function to store images of user-chosen passwords, because it allows for faster adversarial brute-force. In these cases, a key derivation function (KDF), should be used instead. KDFs are functionally similar to hash functions, but are designed in such a way to be uniformly expensive to compute on a broad range of computing devices, for example by requiring not only many CPU cycles but also large amount of memory and high latency. Typical KDFs for cryptographic use are Argon2id [5] and Scrypt [34].
@@
If, in addition to privacy, authenticity of a message is also required, then SKE is not enough, and authenticated encryption (AE) must be used instead. An AE scheme works in a similar way to a SKE, but decryption of a given ciphertext fails if the secret key used for decryption is not the same one used to encrypt the original plaintext. When this happens we write that the decryption procedure returns $\bot$. This allows to check that a ciphertext has not been altered or replaced by a malicious adversary, thereby granting authenticity and integrity of the message. A typical way to implement AE is to append a message authentication code (MAC) to a ciphertext. A MAC is a random-looking bitstring, for example computed through the encrypt-then-MAC procedure with a hash function on a combination of ciphertext and secret key. MACs are useful, among other properties, to check whether a provided key is the correct one to decrypt a ciphertext (without having to actually decrypt the ciphertext first).
@@
As a building block for SKEs, block ciphers are widely used. These are algorithms that typically offer two different interfaces, one for encryption and one for decryption. In encryption mode, they take as input a block of plaintext of fixed bitsize ${S_{\mathsf{B}}}$ (the block size) and an encryption key of ${S_{\mathsf{K}}}$ bits (the key size) and return a block of ciphertext, also of size ${S_{\mathsf{B}}}$. This mapping is undone in decryption mode, provided the same key is used. In other words, block ciphers implement a subset of size at most $2^{S_{\mathsf{K}}}$ of the space of all possible $(2^{S_{\mathsf{B}}})!$ permutations (and their inverses) over ${S_{\mathsf{B}}}$-bit strings. One of the most widely used block ciphers are those from the AES family [20], identified by AES-${S_{\mathsf{K}}}$ (with a block size of 128 bits, and a keysize ${S_{\mathsf{K}}}$ of 128, 192, or 256 bits).
@@ 2.2 Full Disk Encryption
FDE can be implemented using hardware-based or software-based solutions. Hardware-based FDE is typically performed by a dedicated encryption chip, while software-based FDE is achieved using encryption software that runs at the operating system level. Implementation is usually done using standard block ciphers contructions like AES. Some examples of software-based FDE solutions include BitLocker for Windows, FileVault 2 for macOS, and LUKS for Linux [30, 15]. All these implementations have typically only negligible impact on performance compared to a non-FDE system, also thanks to the widespread presence of dedicated CPU instructions to speed up AES computation on personal devices such as smartphones and laptops.
@@ 2.3.2 PD security
Let us first clarify that some constraint is necessary in order to have any hope in the PD game against the adversary. Otherwise, the adversary could submit $O_{0}$ and $O_{1}$ containing completely different (logical) $\mathtt{write}$ accesses to the decoy volumes $V_{0},\ldots,V_{N-2}$, and there would be no way of making the two outcomes indistinguishable, since the adversary holds the passwords to those volumes, so it could trivially verify which of the two patterns was executed. This suggests the need for a minimal rule, stating that, whether $O_{0}$ or $O_{1}$ is executed, the resulting logical contents of the decoy volumes $V_{0},\ldots,V_{N-2}$ must be the same. From the user’s perspective, this basic requirement means that we do not try to disguise the accesses to decoy volumes as something else, both because we do not need to, and because there would be no way of doing it even if we wanted to.
Furthermore, we notice that many PD solutions (including those based on woORAMs) treat $\mathtt{write}$ requests in a completely different way than $\mathtt{read}$ requests, i.e. a $\mathtt{write}$ request could trigger allocation or reshuffling of a certain volume sector. This might happen without breaking the minimal rule described above: if an adversary first $\mathtt{read}$s at a previously unallocated position $\beta_{\mathsf{L}}$, obtaining data $d$, and then $\mathtt{write}$s the same data $d$ at the same location $\beta_{\mathsf{L}}$, this might cause a change of state in the instance without changing its logical content (as mandated by the minimal rule above). This, in turn, would enable an attack where the adversary merely checks whether this state change happened or not. However, we must consider this attack trivial: detecting this kind of state change on a decoy volume would actually not compromise security. Therefore, in order to capture this concept, we also demand that the set of all blocks that are touched by $\mathtt{write}$ requests be the same for both $O_{0}$ and $O_{1}$, for all volumes $V_{0},\ldots V_{N-2}$. We stress that this extra constraint is also very minimal and does not weaken the security guarantees offered in practice; in fact, many previous security definitions demand even stricter constraints, for example by demanding that $O_{0}$ and $O_{1}$ are of the same length, which we do not require here.
@@ 3.1 Design
One big limitation of TrueCrypt is that it can only support a hidden volume if the outer volume is formatted as a FAT filesystem [^5] Latest versions of TrueCrypt on Windows also support NTFS-formatted outer volumes, albeit with serious limitations, among which a waste of 50% volume space.. This is because we need the empty space left by the decoy volume to be contiguous. Most modern filesystems (like ext4) “jump back and forth” on the disk as they are written with data, leaving lots of empty blocks, or “holes”, in the middle. Instead, the FAT filesystem is special, in that it grows incrementally and, up to the physiological holes created by deleted files, occupies all the space up until its last utilised block. This way, one can have the hidden volume start at a certain offset, after the end of the decoy volume, and then follow data allocation linearly. TrueCrypt automatically computes a convenient starting position for the hidden volume (leaving some leeway buffer after the end of the standard volume), and places it among the metadata of the hidden volume’s header. Whereas the hidden volume is assigned a logical size that follows the physical space actually allocated to it, the standard volume is not resized, and keeps logically mapping onto the whole disk. This is crucial in order to not defy deniability: if we resized the standard volume, this information (which leaks the existence of a hidden volume) would be written in the metadata of its file system, which is inspected by the adversary.
@@ 3.2 Operational Model
As already mentioned, once the hidden volume is created, its starting position is final, and it “freezes” the end of the decoy standard volume, limiting the maximum size that it will ever be able to attain. Since the standard volume cannot be resized to accommodate for the hidden volume, and instead keeps mapping onto the whole disk, it is up to the user to not let it grow too much and overwrite the hidden volume. This is achieved both by frequent de-fragmentation of the standard volume, and by actually not writing too much data into it. If one wants to be absolutely safe about data corruption within the hidden volume, the recommendation for the user is to never unlock only the standard volume (except if under coercion), but to always unlock either 1) the hidden volume only, or 2) both hidden and standard volumes, but keeping the latter read-only.
@@ 3.3 Security
In TrueCrypt, once the user surrenders the password of the decoy volume and lets the adversary decrypt it, the only part of the disk contents that remains to be “interpreted” by the adversary is the non-decrypted space in the unused areas of the decoy file system. However, whether this space is actually empty or whether it contains a hidden volume, it will be filled with random bytes that are not readable with the decoy password alone. Therefore, even if a hidden volume is present, the user can plausibly claim that the remaining space is empty and filled with random bytes: the adversary has no way to disprove this claim, or even to question its likelihood, based on the observed disk content.
@@ 3.4 Other Limitations
First of all, as discussed, the standard volume must be formatted as a FAT or NTFS filesystem if a hidden volume is desired. However, FAT is now outdated: it used to be widespread, and it still is on cheap electronics and some data transfer media, but today there is little plausible reason to use it on a personal computing device. Therefore, the mere fact of using FAT raises a red flag. The NTFS option is only available on Windows and comes with other serious limitations.
@@
Another big limitation is the fact that only one hidden volume within each standard volume is supported. This is a problem, because the adversary might reasonably suspect that TrueCrypt is in use exactly to hide something through its PD feature: if we only meant to encrypt a volume, we would reasonably use a more widespread and supported solution like LUKS or BitLocker. A user could claim that they prefer using independent, niche open-source software to secure their data, but the safest course of action is assuming that the adversary might not believe this claim, and will pressure for a second password.
The short answer to this problem is: a robust TrueCrypt-like PD solution should allow simultaneous access to more than two layers of volumes, without revealing how many. That way, we can create a series of volumes with increasingly “private” contents (that could well be all decoys), so as to reveal more than one password to the adversary and convince them, based on the resulting decrypted contents, that we have effectively given up what we were hiding, while in fact we are still holding the password to one more top-secret volume whose existence they have no more reason to suspect.
4 The ‘Legacy’ Shufflecake Scheme
In this section we present (the ‘Legacy’ version of) our Shufflecake scheme, we explain its way of operation, and we provide a security analysis. We do not provide concrete parameters in this section, as we want to present a generic framework that can be adapted to various implementations, but we provide our own choice of parameters in Section 5. In particular, in this section we will treat security as asymptotic, without targeting a concrete bit-security level. We assume a target storage block device of ${S_{\mathsf{D}}}$ blocks, each of them of size ${S_{\mathsf{B}}}$ bits, and a maximum of $N_{\mathsf{MAX}}$ supported volumes per device.
@@ 4.1 Design
At its core, Shufflecake is a block indirection layer on top of an encryption layer. Our indirection layer realizes a mechanism which is already a strong improvement over TrueCrypt, since it fixes two of its crucial limitations: it allows simultaneous use of multiple volumes, and it is filesystem-independent. Decryption keys for every volume are derived by a password and other randomness. Furthermore, the decrypted payload of the header of volume $V_{i}$ with $i>0$ also contains a copy of the header decryption key for volume $V_{i-1}$. This allows to recursively open all volumes present in a device by using a single password, which in turn improves security and user experience, as we will see.
@@
Figure 3: (Legacy) Shufflecake’s disk layout overview.
@@
Figure 4: (Legacy) Shufflecake’s data section layout and slicing.
In Shufflecake ‘Legacy’, we distinguish between logical slices, of size $S_{\mathsf{L}}$ blocks, and physical slices, of size $S_{\mathsf{P}}=S_{\mathsf{L}}+\Delta_{\mathsf{S}}$, where $\Delta_{\mathsf{S}}$ blocks are used to store the encryption IVs for that slice. Logical slices store the data of their respective volumes, while physical slices are used to reserve space on the disk to allocate the encrypted logical slice. A physical slice can either be unallocated (i.e., unclaimed by any volume), or mapped to a single logical slice belonging to some volume $V_{i}$. In the latter case, the mapping from the logical slice index (LSI) $\sigma$ of volume $V_{i}$ to the device’s physical slice index (PSI) $\psi$ is given by a correspondence $\mathsf{SliceMap}_{i}:\sigma\mapsto\psi$ which is just a lookup of $V_{i}$’s slice map (basically, a per-volume array). I/O operations to a logical block address $\beta_{\mathsf{L}}$ of volume $V_{i}$ are performed through the two interfaces $\mathtt{SflcRead}$ and $\mathtt{SflcWrite}$. We will describe later these two interfaces, as well as the structure of the slice map.
@@
Figure 5: (Legacy) Shufflecake’s Device Master Block (DMB).
@@
Figure 6: (Legacy) Shufflecake’s Volume Master Block (VMB).
@@
We now describe more in detail how (‘Legacy’) Shufflecake operates and explain the rationale behind certain design choices. We start by explaining the use encryption on data blocks. Then we look at the indirection layer between physical data on disk and data on the logical volumes, i.e., the mechanism which creates the correspondence between physical and logical slices. We explain how this correspondence is updated when new logical slices are needed as data is written on the volumes. Then, we explain how Shufflecake handles the case of allocation errors, i.e., when a physical slice is mistakenly assigned to more than one volume. Finally, we detail how the instantiation procedure works.
@@
In our ‘Legacy’ case, however, we stick to explicitly random IVs because we want to keep the future possibility to extend Shufflecake with some degree of multi-snapshot security, requiring us to re-encrypt some blocks with a different IV while leaving their content unchanged. For this reason, we use CTR mode instead, and the IV of a block is refreshed at each $\mathtt{SflcWrite}$ for that block, to avoid IV-reuse attacks. This means that all these IVs are stored on-disk.
@@
Concretely, we keep an in-memory, per-device array $\mathsf{prmslices}$ of PSIs of size at most $\mathsf{mxslc}=\lceil\frac{{S_{\mathsf{D}}}-S_{\mathsf{H}}}{S_{\mathsf{P}}}\rceil$ (maximum possible number of slices), and an array of the same size $\mathsf{ofld}$, an “occupation bitfield” telling us which physical slices are occupied (the PSIs are the indexes of the array). Initially, at device instantiation, $\mathsf{prmslices}$ is initialized as $\mathsf{prmslices}[i]:=i$ and then permuted using the efficient Fisher-Yates algorithm [13]. The bitfield $\mathsf{ofld}$ is initialized with all $\mathsf{free}$ values. When a volume is $\mathtt{open}$ed, and we discover the physical slices it maps to, we mark them as $\mathsf{occupied}$ in $\mathsf{ofld}$. The slice allocation algorithm then simply works by repeatedly taking the next element from the pre-shuffled array of PSIs until the bitfield tells us it is free (at which point we take it and mark it as occupied). A stateful occupation counter $\mathsf{octr}$ is kept to facilitate this task, initially set to 00, and then increased up to the first element of $\mathsf{prmslices}$ marked as $\mathsf{free}$ (i.e., the first $\mathsf{octr}$ elements of $\mathsf{prmslices}$ are guaranteed to be $\mathsf{occupied}$ in $\mathsf{ofld}$). This way, when a new slice is required, it is immediate to advance to the next available one. We also need to update $V_{i}$’s slice map $\mathsf{SliceMap}_{i}$ before returning (this is so far the only change that will eventually persist, encrypted, on disk).